# Soft Token Attacks Cannot Reliably Audit Unlearning in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have become increasingly popular. Their emergent capabilities can be attributed to their massive training datasets. However, these datasets often contain undesirable or inappropriate content, e.g., harmful texts, personal information, and copyrighted material. This has promoted research into machine unlearning that aims to remove information from trained models. In particular, approximate unlearning seeks to achieve information removal by strategically editing the model rather than complete model retraining.
Recent work has shown that soft token attacks (STA) can successfully extract purportedly unlearned information from LLMs, thereby exposing limitations in current unlearning methodologies. In this work, we reveal that STAs are an inadequate tool for auditing unlearning. Through systematic evaluation on common unlearning benchmarks (Who Is Harry Potter? and TOFU), we demonstrate that such attacks can elicit any information from the LLM, regardless of (1) the deployed unlearning algorithm, and (2) whether the queried content was originally present in the training corpus. Furthermore, we show that STA with just a few soft tokens ($1-10$) can elicit random strings over 400-characters long. Thus showing that STAs are too powerful, and misrepresent the effectiveness of the unlearning methods.
Our work highlights the need for better evaluation baselines, and more appropriate auditing tools for assessing the effectiveness of unlearning in LLMs.
@@
Haokun Chenthanks: Work done during an internship at Intel Labs. LMU Munich haokun.chen@campus.lmu.de                        Sebastian Szyller Intel Labs contact@sebszyller.com
@@ 1 Introduction
In recent years, large language models (LLMs) have undergone substantial advancements, leading to enhanced performance and widespread adoption. LLMs have demonstrated exceptional performance in various downstream tasks, such as machine translation Zhu et al. (2023), content generation Acharya et al. (2023), and complex problem-solving Xi et al. (2025). Their performance is attributed to their large-scale architectures that require datasets consisting of up to billions of tokens to train effectively Kaplan et al. (2020). These datasets are typically derived from large-scale corpora sourced from public internet text. However, such datasets inadvertently contain harmful or inappropriate content, including instructions for hazardous activities (e.g., bomb-making), violent or explicit material, private information, or copyrighted content unsuitable for applications. Given the sensitive nature of such data, it may be necessary to remove it from the LLM to comply with the local regulations, or internal company policies.
Machine unlearning is a tool for removing information from models Cao and Yang (2015); Bourtoule et al. (2021a). Approximate unlearning usually refers to removing information from models without resorting to retraining them from scratch Zhang et al. (2024a); Eldan and Russinovich (2023a); Izzo et al. (2021), ensuring that the resulting model deviates from a fully retrained version within a bounded error. While numerous studies have proposed various unlearning algorithms, most lack formal guarantees of effectiveness. In fact, prior research has demonstrated that many unlearning techniques can be circumvented through simple rephrasings of the original data Shi et al. (2024). Recent work has shown that a soft token attack (STA) can be used to elicit harmful completions and extract supposedly unlearned information from models Schwinn et al. (2024); Zou et al. (2024).
In this work, we introduce a simple framework for auditing unlearning and demonstrate that STAs are overly powerful, and hence, inappropriate for verifying the effectiveness of approximate unlearning. We show that the auditor can elicit any information from the model, regardless of its training data. Our work highlights the need for better unlearning auditing baselines and methodologies.
We claim the following contributions:
@@
    We introduce a simple auditing framework for unlearning in LLMs (Section 3.2).
@@
The remainder of this paper is organized as follows: In Section 2 we provide an overview of the necessary background, and the related work. Section 3 introduces a general auditing framework, and instantiates it using STA. In Section 4, we demonstrate the efficacy of STA, and subsequently its failure, as a tool for auditing unlearning in LLMs. In Section 5 we discuss additional considerations for auditing unlearning in LLMs. We conclude the paper in Section 6, and highlight some limitations in Section 7.
2 Background & Related work
2.1 Background
Figure 1: Overview of the auditing process using $A_{\textsl{STA}}$. For a perfect unlearning method, $A_{o}$ always correctly audits the model. On the other hand, $A_{\textsl{STA}}$ can elicit the completion regardless of the information in the model – the audit is ineffective.
Large language models (LLMs) process input text through an auto-regressive framework. Given an input sequence of tokens $x_{1:t}$, the model computes the conditional probability distribution $p(x_{t+1}|x_{1:t})$ over the vocabulary at each time-step. The likelihood of the sequence is given by:
$\log p(x_{t+1}|x_{1:t})=\sum_{t=1}^{T}\log p(x_{t}|x_{1:t-1})$ (1)
At inference time, the tokens is generated iteratively by sampling the next token $x_{t+1}$ from this distribution (e.g., via greedy decoding or nucleus sampling Holtzman et al. (2019)), then appending it to the context $x_{1:t}$ for the subsequent step.
Machine unlearning is a tool for removing information from models. Consider a machine learning model $f$ optimized over a training dataset $D_{train}$. When a data owner submits an unlearning request to remove a specified subset $D_{forget}\in D_{train}$, the objective of machine unlearning is to produce an unlearned model $f_{u}$ that eliminates the influence of $D_{forget}$. Machine unlearning methodologies are categorized into two paradigms – exact, and approximate unlearning.
Exact unlearning ensures the output distribution of $f_{u}$ is statistically indistinguishable from that of a model retrained exclusively on the retained dataset $D_{retain}=D_{train}/D_{forget}$. This guarantees provable data removal, satisfying:
$p(f_{u}(x)=y)$ $=p(f_{ret}(x)=y)\quad$ (2)
$s.t.\quad\forall(x,y)\in D_{train},$
where $f_{ret}$ denotes a model retrained from scratch on $D_{retain}$ – which is the most straightforward way of achieving exact unlearning.
The process can be made more efficient by splitting the $D_{train}$ into overlapping chunks, and training an ensemble of models Bourtoule et al. (2021b). During an unlearning request, only the models containing the requested records need to be retrained. For certain classes of models, it is possible to achieve exact unlearning without retraining, e.g. ECO adapts the Cauwenberghs and Poggio (CP) algorithm for exact unlearning within LeNet Huang et al. , and MUSE relabels the target data to achieve unlearning for over-parameterized linear models Yang et al. (2024).
Approximate unlearning, sometimes called in-exact unlearning, relaxes the strict equivalence requirement, instead only requiring that $f_{u}$ approximates $f_{ret}$ within some bounded error. This paradigm relies on empirical metrics or probabilistic frameworks. In LLMs, approximate unlearning is typically accomplished by overwriting the information in the model Eldan and Russinovich (2023a); Wang et al. (2024), guiding the model away from it Feng et al. (2024), or editing the weights and/or activations Liu et al. (2024); Bhaila et al. (2024); Li et al. (2024); Tamirisa et al. (2024); Huu-Tien et al. (2024); Ashuach et al. (2024); Meng et al. (2022a, b).
2.2 Related work
While advances have been made in developing machine unlearning algorithms for LLMs, rigorous methodologies for auditing the efficacy of unlearning remain understudied. Adversarial soft token attacks (STAs) Schwinn et al. (2024) and 5-shot in-context prompting Doshi and Stickland (2024) have been shown to recover unlearned knowledge in models. When model weights can be modified, techniques such as model quantization Zhang et al. (2024e) and retraining on a partially unlearned dataset  Łucki et al. (2024); Hu et al. (2024) have also proven effective in recalling forgotten information.  Lynch et al. (2024) examined eight methods for evaluating LLM unlearning techniques and found that their latent representations remained similar. News and book datasets are used to analyze unlearning algorithms from six different perspectives Shi et al. (2024). It was shown that fine-tuning on unrelated data can restore information unlearned from the LLM Qi et al. (2024), indicating the existing unlearning methods do not actually remove the information but learn a refusal filter instead. Several benchmarks have been developed to evaluate the existing unlearning algorithms. Besides, an unlearning benchmark was introduced based on fictitious author information Maini et al. (2024a). For real-world knowledge unlearning, Real-World Knowledge Unlearning (RWKU) used 200 famous people as unlearning targets Jin et al. (2024), while WDMP focused on unlearning hazardous knowledge in biosecurity, cybersecurity Li et al. (2024).
3 Auditing with Soft Token Attacks
3.1 Adversarial prompts
@@
An adversarial prompt $x_{a}$, is an input prompt to the LLM, obtained by applying the transform $T(\cdot)$ to the base prompt $x_{p}$, $x_{a}=T(x_{p},\textsl{aux})$ in order to elicit a desired completion $c$. $T$ can be any function that swaps, removes or adds tokens; aux denotes any additional needed information. However, such arbitrary attacks are expensive to optimize [^1] Suffix-only attacks allow efficient use of the KV-cache., and difficult to reason about. In practice, $T$ optimizes an adversarial suffix $x_{s}$ that is appended to $x_{p}$ to elicit $c$ Zou et al. (2023). Specifically, we optimize the probability:
@@
In this case, $T$ uses the gradient from the LLM to update $x_{s}$.
3.2 Unlearning auditor
@@
An oracle auditor $A_{o}$ takes an unlearned model $f_{u}$ and the candidate sentences $x_{c}\in X_{c}$ and outputs a ground truth, binary decision $a=\{0,1\}$ indicating whether the given records was part of $D_{train}$ of:
$a=A_{o}(f_{u},X_{c}=D_{forget},\textsl{aux})$ (5)
@@
$A_{o}$ is unrealistic in many scenarios; however, it can be easily instantiated for exact unlearning where $A_{o}$ knows the training data associated with $f$: $aux=\{D_{retain}\}$.
@@
On the other hand, a realistic unlearning $A_{u}$ takes an $f_{u}$, and $D_{forget}$ and outputs a score $s=(0,1)$ indicating whether the records were in $D_{train}$:
$s=A_{u}(f_{u},\{x_{c}\}=D_{forget},\textsl{aux}=\emptyset)$ (6)
@@
$A_{u}$ represents cases where users remove information from models that they did not create, e.g. to prevent harmful outputs.
@@
In this work, we instantiate the soft token attack auditor $A_{\textsl{STA}}$ based on the soft token attacks (STAs) against unlearning Schwinn et al. (2024); Zou et al. (2024). Our auditor compares the relative difficulty of eliciting $c$ for $f_{ft}$ and $f_{u}$. The unlearning procedure is effective if eliciting completions using $f_{u}$ is more difficult than $f_{ft}$.
$s=A_{\textsl{STA}}(f_{u},\{x_{c}\}=D_{forget},\textsl{aux}=\{f_{ft}\}).$ (7)
@@
Figure 1 gives a complete overview of the auditing procedure, and the difference between $A_{0}$ and $A_{\textsl{STA}}$. In Table 1, we summarize the notation.
In the next Section, we show that $A_{\textsl{STA}}$ cannot reliably audit LLMs.
@@ 4.1 Experiment setup
Datasets. To attack unlearning and evaluate its effectiveness, we use two popular benchmark datasets: 1) Who Is Harry Potter?Eldan and Russinovich (2023a), a benchmark that intends to remove information about the world of Harry Potter and the associated characters; WHP hereafter. 2) TOFUMaini et al. (2024a), a dataset of fictional writers that are guaranteed to be absent in the LLM’s training data [^2] This cannot be guaranteed for models published after TOFU..
WHP does not publish a complete dataset. For that reason we use the passages included in the associated Hugging Face page Eldan and Russinovich (2023b). Additionally, we augment it with $20$ $(x_{p}\rightarrow c)$ pairs generated with Llama2-7b-chat-hf. These contain general trivia about the Harry Potter universe.
For TOFU, we use the $10\%$ forget to $90\%$ retain split provided by the authors Maini et al. (2024b).
Models & environment. For all experiments, we use Llama-2-7b-chat-hf Touvron et al. (2023) (Llama2), and Llama-3-8b-instruct Meta (2024) (Llama3) downloaded from Hugging Face. We get the unlearned WHP model from it’s Hugging Face repository Eldan and Russinovich (2023b) (Llama2-WHP).
We implement STA using LLMart Cornelius et al. (2025) – a PyTorch and Hugging Face-based library for crafting adversarial prompts. We use implementations of the unlearning methods from the TOFU Maini et al. (2024c), and NPO Zhang et al. (2024b) repositories. We benchmark the attack against seven different unlearning algorithms: gradient ascent (GA), gradient difference (GDF) Liu et al. (2022), refusal (IDK) Rafailov et al. (2024), knowledge distillation (KL) Hinton (2015), negative preference optimization (NPO) Zhang et al. (2024c), NPO-GDF, NPO-KL.
We run our experiments on a machine equipped with Intel Xeon Gold 5218 CPU, eight NVIDIA A6000, and 256 GB of RAM.
STA soft token attack
$A_{o}$ oracle auditor
$A_{\textsl{STA}}$ STA auditor
$x_{p}$ base prompt (benign)
$x_{s}$ adversarial suffix
$x_{a}$ adversarial prompt ($x_{p}\oplus x_{s}$)
$c$ target completion
$f_{\emptyset}$ base model
$f_{ft}$ fine-tuned model
$f_{u}$ unlearned model
$f_{u-*}$ model unlearned using *
$D_{train}$ training data
$D_{forget}$ forget data
$D_{retain}$ retain data
Table 1: Summary of the notation. ’*’ is replaced with the specific unlearning method.
@@ 4.2 Auditing with attacks
Who Is Harry Potter?. To elicit information about the Harry Potter universe, we initialize the soft tokens using randomly selected hard tokens, and append them to the prompt ($embed(x_{a}=x_{p}\oplus x_{s})$). We then train the soft prompt using AdamW Loshchilov and Hutter (2019) for up to $3000$ iterations; using learning $0.005$, and default $\beta{s}$ – we reiterate that the $x_{p}$ does not change, only the embedded suffix does. If the optimization fails, we double the the number of soft tokens up to the maximum of $16$. We rerun the experiment five times, and report the mean and standard deviation across all prompt and reruns. In Table 2 we report the average number of soft tokens needed to elicit a completion. WHP * denotes the unlearned model with different prompt templates.
Our results show that all information can be generated with $\approx 4-6$ added soft tokens. For all pairs of models, we conduct a t-test under the null hypothesis $\mathcal{H}_{0}$ of equivalent population distributions with $\alpha=0.05$. We use an unpaired Welch’s t-test since sample variances are not equal WELCH (1947). We cannot reject the hypothesis for any of the pairs i.e. $p>0.05$. In other words, for all models, there is not enough evidence to say that eliciting completions is more difficult.
[…]
@@
Table 2: Number of soft tokens required to elicit a completion for a fixed number of iterations. Soft tokens are appended to the prompt. Results are averaged over all prompts in the WHP set and over five runs for each prompt. When we increase the maximum iterations to $10,000$ we can elicit all completions with $1-2$ soft tokens. We do not report the results for Llama2 hf because it was used to generated the data. For comparison we also report the results for Llama3 without any prompt template (N/A).
@@
Table 3: Number of soft tokens required to elicit a completion for a fixed number of iterations. Soft tokens are appended to the prompt. Results are averaged over all prompts in the TOFU set and over five runs for each prompt. When we increase the maximum iterations to $10,000$ we can elicit all completions with $1-2$ soft tokens.
TOFU. For TOFU, we follow the same setup as for WHP– we initialize soft tokens using random hard tokens, and append them; we then train the soft prompt using AdamW for up to $3000$ iterations; using learning $0.005$, and $\beta{s}=(0.9,0.999)$; we double the soft tokens if the optimization fails; we rerun the experiments five times and report the averages across all prompts. In Table 3, we report the number of soft tokens required to elicit the completions. $f_{\emptyset}$ refers to the unmodified baseline model, $f_{ft}$ corresponds to the models fine-tuned on TOFU, followed by the unlearned models.
For all methods, we can elicit the completions with $\approx 3$ appended soft tokens. Similarly to WHP, for all possible pairs of models (within the same architecture), we conduct a t-test under the null hypothesis $\mathcal{H}_{0}$ of equivalent population distributions with $\alpha=0.05$. We use an unpaired Welch’s t-test since sample variances are not equal.
We cannot reject the hypothesis for any of the pairs i.e. $p>0.05$; $f_{u-IDK}$ vs $f_{ft}$ (for Llama2) gives the lowest p-value of $0.509$. In other words, for all models (regardless if trained on TOFU, or used unlearning method), there is not enough evidence to say that eliciting completions is more difficult.
One could argue that used unlearning methods are not effective (when comparing $f_{ft}$ vs $f_{u-*}$), hence they require similar numbers of soft tokens. [^3] Most of these approaches have been in fact been shown ineffective, and susceptible to simple paraphrasing. However, the same holds when compared to $f_{\emptyset}$. In the next section, we demonstrate that the result cannot be attributed to the (in-)effectiveness of the unlearning methods but rather the power of STA.

### v2

@@ Abstract
Large language models (LLMs) are trained using massive datasets. However, these datasets often contain undesirable content, e.g., harmful texts, personal information, and copyrighted material. To address this, machine unlearning aims to remove information from trained models. Recent work has shown that soft token attacks (STA) can successfully extract unlearned information from LLMs. In this work, we show that STAs can be an inadequate tool for auditing unlearning. Using common unlearning benchmarks, i.e., Who Is Harry Potter? and TOFU, we demonstrate that, in a strong auditor setting, such attacks can elicit any information from the LLM, regardless of (1) the deployed unlearning algorithm, and (2) whether the queried content was originally present in the training corpus. Also, we show that STA with just a few soft tokens (1101-10) can elicit random strings over 400-characters long. Thus showing that STAs must be used carefully to effectively audit unlearning. Example code can be found at https://github.com/IntelLabs/LLMart/tree/main
/examples/unlearning
@@
Haokun Chenthanks: Work done while at Intel Labs. Ludwig Maximilian University of Munich Munich Center for Machine Learning (MCML) haokun.chen@campus.lmu.de          Sebastian Szyller11footnotemark: 1 Aalto University contact@sebszyller.com
@@ 1 Introduction
Large language models (LLMs) excel in many downstream tasks, e.g., machine translation Zhu et al. (2023), content generation Acharya et al. (2023), and complex problem-solving Chen et al. (2024). Their performance is attributed to their large-scale architectures that require datasets consisting of up to trillions of tokens to train effectively Kaplan et al. (2020). These datasets are typically derived from large-scale corpora sourced from public internet text. However, such datasets can contain undesirable content, e.g., instructions for building weapons, violent or explicit material, private information, or copyrighted content. Given the sensitive nature of such data, it may be necessary to remove it from the LLM to comply with the local regulations, or internal company policies.
Machine unlearning is a tool for removing information from models Cao and Yang (2015); Bourtoule et al. (2021a). Approximate unlearning usually refers to removing information from models without retraining them from scratch Zhang et al. (2024a); Eldan and Russinovich (2023a); Izzo et al. (2021), ensuring that the resulting model deviates from a fully retrained version within a bounded error. While numerous studies have proposed various unlearning algorithms, most lack formal guarantees. Prior research has demonstrated that many unlearning techniques can be circumvented through rephrasing of the original data Shi et al. (2024). Recent work has shown that a soft token attack (STA) can be used to elicit harmful completions and extract supposedly unlearned information from models Schwinn et al. (2024); Zou et al. (2024).
In this work, we introduce a simple framework for auditing unlearning, and demonstrate that STAs are inappropriate for verifying the effectiveness of approximate unlearning in a strong auditor setting. We show that the auditor can elicit any information from the model, regardless of its training data. We claim the following contributions:
@@
2 Background
@@
Adversarial prompt $x_{a}$ is an input prompt to the LLM, obtained by applying the transform $T(\cdot)$ to the base prompt $x_{p}$: $x_{a}=T(x_{p},\textsl{aux})$ to elicit a desired completion $c$. $T$ can be any function that swaps, removes or adds tokens; aux denotes any additional needed information. Such arbitrary attacks are expensive to optimize, and difficult to reason about. In practice, $T$ optimizes an adversarial suffix $x_{s}$ that is appended to $x_{p}$ to elicit $c$ Zou et al. (2023). These suffix-only attacks also allow efficient use of the KV-cache Pope et al. (2023). Specifically, we optimize the probability:
@@
In this case, $T$ uses the gradient from the LLM to update $x_{s}$. We visualize such attack in Figure 1.
Machine unlearning (MU) aims to remove information from models. Consider a machine learning model $f$ trained using a training dataset $D_{train}$. During an unlearning request to remove a specified subset $D_{forget}\in D_{train}$, the objective of MU is to produce an unlearned model $f_{u}$ that eliminates the influence of $D_{forget}$. There are two types of MU – exact, and approximate unlearning.
Exact unlearning ensures the output distribution of $f_{u}$ is statistically indistinguishable from $f_{ret}$ – a model retrained exclusively on the retained dataset $D_{retain}=D_{train}/D_{forget}$. This guarantees provable data removal, satisfying:
$p(f_{u}(x)=y)$ $=p(f_{ret}(x)=y)\quad$ (3)
$s.t.\quad\forall(x,y)\in D_{train}.$
It can be made more efficient by splitting the $D_{train}$ into overlapping chunks, and training an ensemble Bourtoule et al. (2021b). During an unlearning request, only the models containing the requested records are retrained. For certain classes of models, exact unlearning without retraining is possible, e.g. ECO adapts the Cauwenberghs and Poggio (CP) algorithm for exact unlearning within LeNet Huang et al. , and MUSE relabels the target data to achieve unlearning for over-parameterized linear models Yang et al. (2024).
Approximate unlearning relaxes the strict equivalence requirement, it only requires that $f_{u}$ approximates $f_{ret}$ within some bounded error. It relies on empirical metrics or probabilistic frameworks. In LLMs, approximate unlearning is typically accomplished by overwriting the information in the model Eldan and Russinovich (2023a); Wang et al. (2024), guiding the model away from it Feng et al. (2024), or editing the weights and/or activations Liu et al. (2024); Bhaila et al. (2024); Li et al. (2024); Tamirisa et al. (2024); Huu-Tien et al. (2024); Ashuach et al. (2024); Meng et al. (2022a, b).
Figure 1: STA combines $x_{p}$ with the optimized $x_{s}$.
3 Related work
While advances have been made in developing machine unlearning algorithms for LLMs, rigorous methodologies for auditing the efficacy of unlearning remain understudied. Adversarial soft token attacks (STAs) Schwinn et al. (2024) and 5-shot in-context prompting Doshi and Stickland (2024) have been shown to recover unlearned knowledge in models. When model weights can be modified, techniques such as model quantization Zhang et al. (2024e) and retraining on a partially unlearned dataset  Łucki et al. (2024); Hu et al. (2024); Chen et al. (2025) have also proven effective in recalling forgotten information.  Lynch et al. (2024) examined eight methods for evaluating LLM unlearning techniques and found that their latent representations remained similar. News and book datasets are used to analyze unlearning algorithms from six different perspectives Shi et al. (2024). It was shown that fine-tuning on unrelated data can restore information unlearned from the LLM Qi et al. (2024), indicating the existing unlearning methods do not actually remove the information but learn a refusal filter instead. Several benchmarks have been developed to evaluate the existing unlearning algorithms. Besides, an unlearning benchmark was introduced based on fictitious author information Maini et al. (2024a). For real-world knowledge unlearning, Real-World Knowledge Unlearning (RWKU) used 200 famous people as unlearning targets Jin et al. (2024), while WDMP focused on unlearning hazardous knowledge in biosecurity, cybersecurity Li et al. (2024).
4 Auditing with Soft Token Attacks
@@
An oracle auditor $A_{o}$ takes an unlearned model $f_{u}$ and the candidate sentences $x_{c}\in X_{c}$, and outputs a ground truth, binary decision $a=\{0,1\}$ indicating whether the given records was part of $D_{train}$ of:
$a=A_{o}(f_{u},X_{c}=D_{forget},\textsl{aux}).$ (4)
@@
$A_{o}$ is unrealistic but it can be easily instantiated for exact unlearning where $A_{o}$ knows the training data associated with $f$: $aux=\{D_{retain}\}$.
@@
A realistic unlearning auditor $A_{u}$ takes an $f_{u}$, and $D_{forget}$ and outputs a score $s=(0,1)$ indicating whether the records were in $D_{train}$:
$s=A_{u}(f_{u},X_{c}=D_{forget},\textsl{aux}=\emptyset)$ (5)
@@
$A_{u}$ represents cases where users unlearn facts from models that they did not create, e.g. to prevent harmful outputs.
@@
In this work, we instantiate the soft token attack auditor $A_{\textsl{STA}}$ based on the soft token attacks (STAs) against unlearning Schwinn et al. (2024); Zou et al. (2024). $A_{\textsl{STA}}$ compares the relative difficulty of eliciting $c$ for $f_{ft}$ and $f_{u}$. The unlearning procedure is effective if eliciting completions using $f_{u}$ is more difficult than $f_{ft}$.
$s=A_{\textsl{STA}}(f_{u},X_{c}=D_{forget},\textsl{aux}=\{f_{ft}\}).$ (6)
Crucially, Schwinn et al., 2024 optimized $x_{s}$ only w.r.t. the affirmative beginning of $x_{c}$ ($x_{p}=$Who is Harry Potter?’, $x_{c}=$Harry Potter is’). Contrary to that minimal setting, our strong auditor $A_{\textsl{STA}}$ optimizes $x_{s}$ w.r.t. the entire $x_{c}$. Our goal is to study the extreme scenario where the auditor expects the exact completions for, e.g. to check for copyrighted content, or personal information.
@@
In the Appendix, Figure 3 gives an overview of the auditing procedure, and the difference between $A_{o}$ and $A_{\textsl{STA}}$; Table 2 summarizes the notation.
@@ 5.1 Experiment setup
Datasets. For evaluation, we use two popular benchmark datasets: 1) Who Is Harry Potter?Eldan and Russinovich (2023a)(WHP) that intends to remove formation about the world of Harry Potter. WHP does not publish a full dataset. Hence, we use the snippets from the associated Hugging Face page, and we augment them with 2020 $(x_{p}\rightarrow c)$ Harry Potter trivia pairs generated with Llama2. 2) TOFUMaini et al. (2024a)is a dataset of fictional writers, designed to be absent from the LLM’s training data. Note that models released after TOFU was published might contain its records. We use the provided $10\%$ forget to $90\%$ retain split Maini et al. (2024b).
Models. We use Llama-2-7b-chat-hf Touvron et al. (2023) (Llama2), and Llama-3-8b-instruct Meta (2024) (Llama3). We get the unlearned WHP model from Hugging Face Eldan and Russinovich (2023b) (Llama2-WHP).
Implementation. We implement STA using LLMart Cornelius et al. (2025) – a PyTorch-based library for crafting adversarial prompts. We use implementations of the unlearning methods from the TOFU Maini et al. (2024c), and NPO Zhang et al. (2024b) repositories. We benchmark the attack against seven different unlearning algorithms: gradient ascent (GA), gradient difference (GDF) Liu et al. (2022), refusal (IDK) Rafailov et al. (2024), knowledge distillation (KL) Hinton (2015), negative preference optimization (NPO) Zhang et al. (2024c), NPO-GDF, NPO-KL.
@@ 5.2 Auditing with attacks
Who Is Harry Potter?. To elicit completions, we initialize the soft tokens using randomly selected hard tokens, and append them to the prompt $x_{a}=embed(x_{p})\oplus embed(x_{s})$. We then train the soft prompt using AdamW Loshchilov and Hutter (2019) for up to $3000$ iterations; using $lr=0.005$, and $\beta{s}=(0.9,0.999)$. $x_{p}$ does not change, only the embedded suffix does. If the optimization fails, we double the number of soft tokens up to the maximum of $16$. We report the mean and standard deviation over five independent runs across all prompts. In Table 1(a) we report the average number of soft tokens needed to elicit a completion. WHP * denotes the unlearned model with different prompt templates.
We show that all target completions can be generated with $\approx 4-6$ added soft tokens. For all pairs of models, we conduct a t-test under the null hypothesis $\mathcal{H}_{0}$ of equivalent population distributions with $\alpha=0.05$. We use an unpaired Welch’s t-test since sample variances are not equal WELCH (1947). We cannot reject the hypothesis for any of the pairs i.e. $p>0.05$. In other words, for all models, no significant evidence that eliciting completions is more difficult.
@@
(a) WHP results with different prompt templates.
@@
(b) TOFU results with different unlearning methods.
Table 1: Number of soft tokens needed to elicit a completion for a fixed number of iterations; averaged over all prompts in each set and over five runs per prompt. When increasing the maximum iterations to 10,00010{,}000, all completions can be elicited with $1\text{--}2$ soft tokens.
TOFU. We follow the same setup as for WHP. In Table 1(b), we report the number of soft tokens required to elicit the completions. $f_{\emptyset}$ refers to the unmodified baseline model, $f_{ft}$ are the models fine-tuned on TOFU, followed by the unlearned models.
For all methods, we can elicit the completions with $\approx 3$ soft tokens. Similarly to WHP, for all pairs of models (within the same architecture), we conduct Welch’s t-test. We cannot reject the hypothesis for any of the pairs i.e. $p>0.05$; $f_{u-IDK}$ vs $f_{ft}$ (for Llama2) gives the lowest p-value of $0.509$. For all models, there is not enough evidence to say that eliciting completions is more difficult.
One could argue that the unlearning methods used are not effective (when comparing $f_{ft}$ vs $f_{u-*}$), hence they require similar numbers of soft tokens. In fact, most of these approaches have already been shown to be ineffective and susceptible to simple paraphrasing Shi et al. (2024). However, the same holds when compared to $f_{\emptyset}$. In the next section, we demonstrate that the result cannot be attributed to the (in-)effectiveness of the unlearning methods but rather the power of STA.